OpenStudy (anonymous):

algerbra 1..questions and anwsers..help please. If the sequence is geometric, what is the common ratio? 500, 100, 20, 4,... 5 125 1/5 not geometric Which of the following is the formula for the geometric sequence? 2, 6, 18, 54, ... an = 3(2)n-1 an = 2(4)n-1 an = 2(3)n-1 Which of the following is the formula for the geometric sequence? 320, 80, 20, 5,... an = 320(1/4)n-1 an = 320(4)n-1 an = 4(320)n-1

OpenStudy (anonymous):

1. 1/5

OpenStudy (anonymous):

first 500/100=100/20=20/4=5 second an = 2(3)n-1 third an = 320(1/4)n-1

OpenStudy (anonymous):

oh first is 1/5